# Scaling Laws for Black-box Adversarial Attacks

### v1

@@ 4.3 Scaling Laws on Vision Language Models
  • $\alpha$ varies significantly across models, revealing the considerable differences in robustness between VLMs.
  • Since VLMs are trained on multimodal data, different adversarial targets significantly influence the attack success rates. This variability may stem from the inherent characteristics of the target labels, as naturally some labels (such as “deer”) are less representative compared to labels like “cat”. This phenomenon, often referred to as class imbalance, has been discussed in prior studies [44].
When compared to image classifiers, VLMs demonstrate superior robustness across all chosen adversarial targets. This supports the general opinion that multimodal training leads to more robust representations [40], enabling the model to focus more on the essential features of the image rather than specific attack patterns. Furthermore, in contrast to traditional views that scaling model parameters results in increased robustness, we find that among the three target models, Phi-3-vision exhibits the highest robustness. This observation suggests that the scaling laws derived in [8] may not be directly applicable to multimodal models.
@@ 5 Extending Scaling Laws to Vision Encoders
In this section we extend our findings of scaling laws from image classifiers to multimodal foundation models such as CLIPs. Sec. 5.1 shows the extension of Eq. (2) to image embeddings, while Sec. 5.2 demonstrates high success rates over commercial multimodal large language models (MLLMs) under our framework.
Figure 5: Demonstrations of scaling laws when we use CLIPs as surrogate models under $\epsilon=8/255$. We take ASR measured by Metric 1 to ensure our results are of high confidence.
@@ 5.1 Attacking Image Embeddings
Note that besides an empirical loss function, Eq. (11) can be also explained as minimizing the distance between the embeddings of the adversarial examples and the target images on a low-dimension embedding manifold.
5.2 Experiments and Evaluations
@@
Surrogate Models. We select 12 pretrained open-source CLIP models via OpenCLIP [15]. In contrast to previous works that typically use fewer than 4 surrogate models, our ensemble choice is sufficient to yield robust results. These CLIP models vary in architecture, hyperparameters, input image sizes, and even training datasets.
Black-box Target Models. We choose 5 commercial MLLMs from the MultiTrust leaderboard with high robustness as our target black-box models: GPT-4o, GPT-4-turbo [66], Claude-3.5-sonnet https://www.anthropic.com/news/claude-3-family, Hunyuan [75], and Qwen [6].
Evaluation Metrics. We employ two types of evaluation metrics in this study. The first metric is the same as in Sec. 4.3. The second metric focuses on the features of adversarial examples, whose goal is to evaluate the insertion of features within adversarial targets into these examples. It requires the test model to generate a description and asks GPT whether the description mentions the adversarial target. Both metrics take ASR as the final score, which we refer to as Metric 1 and Metric 2 in the following sections for simplicity. Test prompts can be found in Appendix C.
ASR(%) Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
@@
We also conduct extensive experiments using Metric 2. Table 2 shows our attack success rates across some of the test LLMs (results of other models can be found in Appendix D). Remarkably, under $\epsilon=16/255$, our CLIP ensemble achieves over 90% success on the GPT series and more than 60% on other commercial LLMs. These results highlight the inherent vulnerabilities of commercial LLMs and underscore the need for the development of more robust and trustworthy multimodal foundation models.
6 Further Studies
(a) Visualization of adversarial perturbations.
(b) Visualization of saliency maps on different models.
Figure 6: Visualization for analysis. (a) It demonstrates the trend of the increasing semantic interpretability of adversarial perturbations over different target objects. (b) It shows the saliency maps of normal image classifiers and adversarially trained models, measured by the first step of the gradient descent iterations.
In this section, we conduct further studies of our scaling laws derived in the past sections. We attempt to explain the existence of scaling laws from two perspectives: first through the visualization of adversarial perturbations, and second by showing that scaling laws do not hold in adversarially trained models.
@@ 6.1 Visualizations of adversarial perturbations
As shown in Fig. 6(a), the semantic meaning of our adversarial perturbations also increases as the cardinality of the model ensemble grows. This suggests that the attack success rates are closely related to the interpretability of the perturbations, aligning with human observations and preferences [18]. Since the ultimate goal of adversarial attacks is to generate high-quality adversarial examples, it is preferable to have interpretable perturbations, as this reinforces the idea that the scaling effectively helps learn common features, rather than merely converging to local optima.
@@
When we enlarge $\epsilon$ from $8/255$ to $16/255$ or even larger, the affect of the natural image goes down to 0, and the adversarial attack task degenerates to an image generation task. We claim that our method of scaling model ensemble still remains good performance under the degenerated image generation task, where the visualized results are shown in Appendix E.
6.2 Limitations of Scaling Laws
Though we have derived scaling laws from normal image classifiers, the results on adversarially trained models disapprove the fact of scaling in our experiments. The experiment settings and results can be found in Appendix F, where we demonstrate that enlarging the cardinality of model ensemble do not actually increase transfer attack success rates over all these adversarial defense models. By visualization of saliency maps between adversarially trained models and normal image classifiers in Fig. 6(b), we observe that they have a completely different focus when performing gradient descent on $\bm{x}$. The result supports that scaling laws exist only when the target models and surrogate models follow the same distribution (i.e., are independently sampled from an identical model set), which corroborates the i.i.d. assumption in Theorem 1.
@@ 7 Conclusion
In this work, we investigate the scaling laws for black-box adversarial attacks. Based on state-of-the-art transfer attack algorithms, we first provide theoretical motivations on the convergence of adversarial examples and training loss, and then propose the scaling laws empirically by our observations that the attack success rate (ASR) and the loss are converging in an exponential manner. We conduct extensive experiments over both image classifiers and modern vision language models (VLMs) to state our scaling laws are robust enough under different scenarios. Leveraging our insights of scaling laws, we emphasize possible practical applications by attacking commercial Multimodal Large Language Models (MLLMs) targeting on an ensemble of CLIP vision encoders. We eventually analyze the reasons of the scaling laws for adversarial attack, from the perspectives of perturbation visualization and evaluation on adversarially trained models. We believe the findings in this work can help deepen the understanding of the weaknesses of vision models regarding adversarial robustness in a black-box context, ultimately contributing to the development of more robust and trustworthy multimodal foundation models.
@@ Appendix B Experiment Settings for Scaling Laws
adv-inception-v3 adversarially trained
ens-adv-inception-resnet-v2 adversarially trained
@@
Model name Architecture Settings
ViT vit-b-16 default
vit-b-32 default
vit-l-16 default
vit-l-32 default
vit-b-16 finetuned on Imagefolder
@@
BeiT beit-B-16 trained on ImageNet 21k
@@
Appendix C Test Prompts as Metrics
@@
Appendix D Experiment Results for MLLMs
@@
ASR(%) Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
Gemini-1.5 Pro 0.4 0.92 0.16 0.50 0.35 0.82 0.29 0.82 0.38 0.83
Hunyuan-Vision 0.58 0.95 0.23 0.53 0.59 0.90 0.42 0.88 0.52 0.95
Qwen-VL Plus 0.75 0.98 0.28 0.54 0.65 0.91 0.52 0.92 0.69 0.97
Llama-3.2 90B 0.48 0.88 0.21 0.48 0.42 0.87 0.39 0.87 0.4 0.86
Pixtral Large 0.49 0.88 0.19 0.49 0.46 0.87 0.36 0.77 0.44 0.87
Deepseek-VL 7B 0.54 0.96 0.17 0.52 0.52 0.87 0.44 0.91 0.51 0.93
Table 6: ASR under Metric 2 measured across different settings of $\epsilon$. Following Table 2, we extend our evaluations to most popular models (both commercial and open-sourced) to obtain more generalized results.
@@
Figure 9: Results on attacking adversarially trained models in the untargeted metric setting. No obvious scaling trends can be observed from the results, and the attack success rate remains consistent regardless of ensemble cardinality.
In Table 5 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [71], OpenCLIP [15], and Timm [87]. We select 6 commercial MLLMs as our black-box test models - GPT-4o, GPT-4 Turbo [66], Claude-3.5 Sonnet https://www.anthropic.com/news/claude-3-family, Gemini-1.5 Pro §§§https://deepmind.google/technologies/gemini/pro/, Hunyuan [75], Qwen [6] - and 3 open-source MLLMs - LLama-3.2 90B https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Pixtral Large https://mistral.ai/news/pixtral-large/, and DeepSeek-VL 7B [19]. In Table 2 we only list part of the results for better demonstration of our capability in creating powerful adversarial examples, therefore in Table 6 we list experiment results for rest of these models. From the results we observe that the high attack success rates of our adversarial examples are generalized results, and the transferability of our adversarial examples is strong enough to fool most popular MLLMs. Compared to previous methods that tried to attack commercial MLLMs by visual adversarial examples [27], our methods exhibit a 40% higher success rate. This demonstrates the strength of our core inspiration of scaling, and highlights the practical applications of our methods. Meanwhile, we find the recently released open-source models Llama-3.2 90B and Pixtral Large exhibiting a better performance than other models generally, which indicates that the AI community is running towards developing more robust, safe and reliable MLLMs.
Appendix E Degenerated Image Generation Task
@@
Appendix F Results on Adversarially Trained Models
We provide our experiment settings and results for Sec. 6.2. We still adapt the adversarial examples generated from normal image classifiers and test them on adversarially trained models, as in Sec. 4.3.
Black-box Test Models: We select 9 pretrained models from robustbench [16] leading board with different architectures and defense methods - architectures include ResNet [35], Swin-Transformers [59], ConvNeXt [60], ViT [28] and Wide-ResNet [98] , while defense methods include Mean-Centered feature sparsification [2], designing better benchmarks [55], combining with ConvStem architecture [74], incorporating with generalizable design principles [69], data filtering [12], randomly masking gradients [65], and using FGSM optimizer during adversarial training [89]. By selecting these models as test subjects, we ensure that a sufficient variety of defense mechanisms are represented, allowing us to draw reliable conclusions.
Evaluation Metrics: Since adversarially trained models typically incorporate defense strategies against adversarial examples, we adjust our evaluation approach. Specifically, we measure the attack success rate based on whether the test models classify the adversarial examples differently from the natural images, a widely used method in untargeted attack evaluations.
As we observe from Fig. 9, a larger model ensemble does not indicate a higher transfer attack success rate. Instead, under our untargeted attack metric, the attack success rate remains consistent across all adversarially trained models. This reveals the fact that scaling laws do not hold over adversarially trained models. By visualizations in Sec 6.2, we claim this is because the gap between normal image classifiers and adversarially trained classifiers on image understanding. Adversarially trained classifiers have a completely different focus on image features, making them insensitive to attack patterns crafted from normal image classifiers. The results show scaling laws only exist when models are sampled from similar data distribution (i.e. normal datasets that are sampled from physical world or training methods that do not exclude specific features of the data), which supports the assumption in Sec. 4.1 of i.i.d. models. Meanwhile, though the attack success rate only slightly changes with ensemble cardinality, we observe from Fig. 9 that many adversarially trained models have bad classification accuracy. This is because adversarially trained models try to strike a balance between accuracy and robustness [68], therefore losing accuracy under small noise.

### v2

@@ 3.4 Scaling Laws on MLLMs
  • $\alpha$ varies significantly across models, revealing the considerable differences in robustness between MLLMs.
  • Since MLLMs are trained on multimodal data, different adversarial targets significantly influence the attack success rates. This variability may stem from the inherent characteristics of the target labels, as naturally some labels (such as “deer”) are less representative compared to labels like “cat”. This phenomenon, often referred to as class imbalance, has been discussed in prior studies [46].
When compared to image classifiers, MLLMs demonstrate superior robustness across all chosen adversarial targets. This supports the general opinion that multimodal training leads to more robust representations [42], enabling the model to focus more on the essential features of the images rather than specific attack patterns. Furthermore, in contrast to traditional views that scaling model parameters results in increased robustness, we find that among the three target models, Phi-3-vision exhibits the highest robustness. This suggests that the scaling laws derived previously in [5] may not be directly applicable to multimodal models.
@@ 4 Extending Scaling Laws to Vision Encoders
Figure 6: Scaling Laws when we use CLIPs as surrogate models under $\epsilon=8/255$. We take ASR measured by Metric 1 to ensure our results are of high confidence.
In this section, we extend our findings of scaling laws from image classifiers to multimodal foundation models such as CLIP [70]. Sec. 4.1 shows the extension of Eq. (2) to image embeddings, while Sec. 4.2 demonstrates high success rates over commercial multimodal large language models (MLLMs) under our framework.
@@ 4.1 Attacking Image Embeddings
Note that besides an empirical loss function, Eq. (10) can also be explained as minimizing the distance between the embeddings of the adversarial examples and the target images on a low-dimension embedding manifold.
4.2 Experiments
@@
Surrogate Models. We select 12 pretrained open-source CLIP models via OpenCLIP [12]. In contrast to previous works that typically use fewer than 4 surrogate models, our ensemble choice is sufficient to yield robust results. The full list of our surrogate clips can be found in Appendix. E.
Target Models. We select 6 commercial MLLMs as our black-box test models - GPT-4o, GPT-4 Turbo [68], Claude-3.5 Sonnet  [^2] https://www.anthropic.com/news/claude-3-family, Gemini-1.5 Pro  [^3] https://deepmind.google/technologies/gemini/pro/, Hunyuan [30], Qwen [28] - and 3 open-source MLLMs - LLama-3.2 90B  [^4] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Pixtral Large  [^5] https://mistral.ai/news/pixtral-large/, and DeepSeek-VL 7B [17].
ASR Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
@@
Gemini-1.5 Pro 0.4 0.92 0.16 0.50 0.35 0.82 0.29 0.82 0.38 0.83
Hunyuan-Vision 0.58 0.95 0.23 0.53 0.59 0.90 0.42 0.88 0.52 0.95
Qwen-VL Plus 0.75 0.98 0.28 0.54 0.65 0.91 0.52 0.92 0.69 0.97
Llama-3.2 90B 0.48 0.88 0.21 0.48 0.42 0.87 0.39 0.87 0.4 0.86
Pixtral Large 0.49 0.88 0.19 0.49 0.46 0.87 0.36 0.77 0.44 0.87
Deepseek-VL 7B 0.54 0.96 0.17 0.52 0.52 0.87 0.44 0.91 0.51 0.93
@@
Evaluation Metrics. We employ two types of evaluation metrics in this study. The first metric is the same as in Sec. 3.4. The second metric focuses on the features of adversarial examples, whose goal is to evaluate the insertion of features within adversarial targets into these examples. It requires the target model to generate a description and asks GPT whether the description mentions the adversarial target. Both metrics take ASR as the final score, which we refer to as Metric 1 and Metric 2 in the following sections for simplicity. Test prompts can be found in Appendix D.
@@
We also conduct extensive experiments under Metric 2. From Table. 2 we observe that the high attack success rates of our adversarial examples are generalized results, and the transferability of our adversarial examples is strong enough to fool most popular MLLMs. Compared to previous methods that tried to attack commercial MLLMs by visual adversarial examples [26], our methods exhibit a 40% higher success rate. This demonstrates the strength of our core inspiration of scaling, and highlights the practical applications of our methods. Meanwhile, we find the recently released open-source models Llama-3.2 90B and Pixtral Large exhibiting a better performance than other models generally, which indicates that the AI community is running towards developing more robust, safe and reliable MLLMs.
5 Discussions
In this section, we further demonstrate the scaling laws by using different optimizers, visualizing the adversarial perturbations, and collecting the fitting error.
5.1 Scaling Laws are General Principles
The idea that enlarging the model ensemble induces higher attack success rates has been proposed in previous work [94]. Nevertheless, while prior work usually provides qualitative guesses or bounds, we exploit linear scaling laws by extensive quantitative experiments. While our scaling laws work well with targeted attack settings and the SSA-CWA algorithm, further experiment results reveal that our scaling laws are general principles.
(a) Under untargeted setting.
(b) Across different algorithms.
Figure 7: Ablations on untargeted settings and other algorithms. We mainly follow the experiment configurations in Sec. 3.1. For untargeted attacks, we set $\epsilon=2/255$. For other algorithms, we leverage cat as the target class and plot the average ASR over all target models. (a) It demonstrates our linear scaling laws work under untargeted settings. (b) It supports our claim that linear trends occur on variance-reduction algorithms, while MI-FGSM exhibits inferior performance due to gradient conflicts.
Untargeted Attack Setting. Untargeted attacks are similar to targeted attacks except for a different training objective from Eq. (2), and all our algorithms are obviously compatible with the untargeted setting:
$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (11)
Our experiment results in Fig. 7(a) serve as a clear sign that scaling laws remain valid over untargeted settings. We mainly follow our settings in Sec. 3.1, while we set $\epsilon=2/255$, and measure the results on 3 target models. Taking an extremely small $\epsilon$ is necessary for demonstrating our linear scaling laws, otherwise the attack success rate will easily reach $100\%$ and therefore come to a plateau. The results validate our claim that scaling laws exist regardless of specific attack targets.
Across Different Attack Algorithms. Besides SSA-CWA, we also provide experimental validations of different attack algorithms in Fig. 7(b). As stated in Sec. 2.2, severe gradient conflicts can be observed when we iteratively update the input image $\bm{x}$ if we directly perform backward propagation over the sum of output logits (e.g., as in MI-FGSM). Gradient conflicts prevent the input $\bm{x}$ from being updated properly and lead to missing information of model weaknesses. However, advanced algorithms, such as SVRE [92] and CWA [9] address this issue by leveraging optimizers that can reduce the variance of gradients. Our experiment results support our claim that scaling laws occur only on such algorithms. Note that the attack success rate of MI-FGSM also becomes higher with increasing ensemble cardinality, further validating that our theoretical insight from Theorem 2 is regardless of specific attack methods.
@@ 5.2 Visualizations of Adversarial Perturbations
Figure 8: Visualization of adversarial perturbations. A clear trend of the increasing semantic interpretability of adversarial perturbations over different target objects.
As shown in Fig. 8, the semantic meaning of our adversarial perturbations also increases as the cardinality of the model ensemble grows. This suggests that the attack success rates are closely related to the interpretability of the perturbations, aligning with human observations and preferences [15]. Since the ultimate goal of adversarial attacks is to generate high-quality adversarial examples, it is preferable to have interpretable perturbations, as this reinforces the idea that the scaling effectively helps learn common features rather than merely converging to local optima.
@@
When we enlarge $\epsilon$ from $8/255$ to $16/255$ or even larger, the effect of the natural image goes down to 0, and the adversarial attack task degenerates to an image generation task. We claim that our method of scaling model ensemble still performs well under the degenerated image generation task, where the visualized results are shown in Appendix F.
5.3 Fitting Error of Scaling Law Curve
To comprehensively demonstrate the universality of our scaling laws, we provide the MSE fitting error from Fig. 3 in Table.3. It can be observed that the small fitting error acts as a quantitative validation of our scaling laws. Thus, our scaling laws remain a scalable way to enhance the transferability of adversarial examples.
Targets Cat Car Ship Deer Bird
Metric ASR Loss ASR Loss ASR Loss ASR Loss ASR Loss
Fitting Error 0.0040 0.34 0.0040 0.17 0.0038 0.46 0.0025 0.05 0.0023 0.27
Table 3: The averaged MSE fitting error of our curves in Fig. 3. Note that ASR usually has range $[0,1]$, while loss has range $[0,10]$. The fitting error is averaged on all test models, not only the test models that are fitted with lines for demonstration.
@@ 6 Conclusion
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models like AT and DiffPure (Sec. 3.3), and modern MLLMs (Sec. 3.4 and Sec. 4) using various optimizers (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer attacks.
Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than focusing on designing intricate algorithms on a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy multimodal large foundation models.
@@ Appendix B Experiment Settings for Scaling Laws
ViT vit-b-16 default
vit-b-32 default
vit-l-16 default
vit-l-32 default
vit-b-16 pretrained on ImageNet-21K
Model name Architecture Settings
@@
BeiT beit-B-16 pretrained on ImageNet-21k
@@
Twins-SVT twins-svt-large default
NFNet nfnet-l0 default
@@
Appendix C SSA-CWA Algorithm Declaration
We provide the pseudo-code version of SSA-CWA implementation in Algorithm 1. Their combination is induced from  [63] and  [9], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their original papers.
Algorithm 1 SSA-CWA Algorithm
1:Input: natural image $\bm{x}_{\text{nat}}$, label $\bm{y}$, $L_{\infty}$ constraint $\epsilon$, iterations $T$, loss function $\mathcal{L}$, spectrum transformation function $\mathcal{T}$, model ensemble $\{f_{i}\}_{i=1}^{N}$, spectrum transformation times $M$, decay factor $\mu$, step sizes $\alpha,\beta$.
2:Output: An adversarial example $\bm{x}$
3:for t = 0 to T-1 do
4:     Update $\bm{x_{t}}$ by $\bm{x}^{0}_{t}=x_{t-1}$
5:     for i = 1 todo
6:         for j = 1 todo $\triangleright$ Spectrum transformation
7:              Obtain transformed image $\mathcal{T}(\bm{x}^{j}_{t,i})$
8:              Compute $\bm{g}_{i}^{j}=\nabla_{\bm{x}^{j}_{t,i}}\mathcal{L}(\mathcal{T}(\bm{x}^{j}_{t,i}),\bm{y})$
9:         end for
10:         Compute gradient $\bm{g}_{i}=\frac{1}{M}\sum_{j=1}^{M}\bm{g}^{j}_{i}$
11:         Update inner momentum by $\bm{m}=\mu\cdot\bm{m}+\frac{\bm{g}_{i}}{\left\lVert\bm{g}_{i}\right\rVert_{2}}$
12:         Update $\bm{x}^{i}_{t}=\text{clip}(\bm{x}_{\text{nat}},\epsilon)\{\bm{x}^{i-1}_{t}+\beta\cdot\bm{m}\}$
13:     end for
14:     Compute gradient $\bm{g}=\bm{x}_{t}^{N}-\bm{x}_{t}^{0}$
15:     Update outer momentum by $\bm{h}=\mu\cdot\bm{h}+\frac{\bm{g}}{\left\lVert\bm{g}\right\rVert_{2}}$
16:     $\bm{x}_{t}=\text{clip}(\bm{x}_{\text{nat}},\epsilon)\{\bm{x}_{t}^{N}+\alpha\cdot\text{sign}(\bm{h})\}$
17:end for
18:Return: $\bm{x}_{T-1}$
Appendix D Test Prompts as Metrics
@@
Appendix E Experiment Settings for Attacking MLLMs
@@
In Table. 6 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [70], OpenCLIP [12], and Timm [84].
Appendix F Degenerated Image Generation Task
@@
Appendix G Related Work
G.1 Black-Box Attacks
G.2 Ensemble Attacks
Prior studies [41] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization. Recent research [94] has theoretically defined transferability error and sought to minimize it through model ensemble. Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [59].
G.3 Scaling Laws
In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [50, 5]. However, the broader concept of scaling remains underexplored. While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [94], these effects have not been systematically analyzed. Recent work Wei et al. [83] also explore the scaling law of attack success rate when jailbreaking large language models [99, 10] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models. Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.